[Hallc_running] [New Logentry] RC daily update

hszumila at jlab.org hszumila at jlab.org
Tue Sep 25 08:45:01 EDT 2018

Logentry Text:
We got some beam in the Hall last night. The harp scans upstream looked ok, but the scans at the target looked too big: https://logbooks.jlab.org/entry/3597193
Jay advised that we could proceed with the big BPM calibration. This for some reason was failed/terminated: https://logbooks.jlab.org/entry/3597251
Jay and crew requested we begin the the deflection test. Then nothing else happened until the end of shift. 

Jay found this morning that the Moller quad had been turned on during beam tuning last night and is likely the culprit for the large beam at our target. https://logbooks.jlab.org/entry/3597276 We will re-swipe as soon as beam comes back this morning. 

Today the plan is to push to high currents for Hall C. This is the priority over any other hall receiving beam and will begin after some BLA work in the late morning.

We will meet today in the second floor room of the counting house at 4:30pm (note the change in time!). Connection details:


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3597328
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20180925/4bbd1a87/attachment.html>

More information about the Hallc_running mailing list